# LLM+KG’24: The International Workshop on Data Management Opportunities in Unifying Large Language Models + Knowledge Graphs, August 25, 2024, Guangzhou, China

### v1

@@ 4.1 KG Question Answering
Multi-Hop Question Generation: The goal of multi-hop question generation is to formulate questions demanding advanced reasoning over multiple sentences, including their associated answers [66].li2023multi created a KG Enhanced Language Model (KGEL), aiming to replicate human-like reasoning for multi-hop questions. Their method for generating multi-hop questions reflects the way humans approach complex questioning. Their strategy encompasses three key phases: (1) Gaining an understanding of the context by employing a pre-trained GPT-2 language model, (2) Merging information and reasoning through a sophisticated KG and an answer-aware dynamic Graph Attention Network (GAT), and (3) Executing the question generation using a multi-head attention generation module that capitalizes on improved latent representations. [67] research primary emphasis revolved around formulating inquiries by employing KGs in conjunction with the T5 language model. They harnessed the language model’s capabilities to generate questions while also applying a technique to obstruct self-attention in the encoder, all to train the model with a deliberate emphasis on maintaining the inherent structure of the graph. Their research didn’t target multi-hop question answering.
@@
Complex or Multi-hop Question Answering: Complex question answering over KGs involves multiple subjects, conveys interconnected relations, involves numerical operations [68], and demands multi-hop reasoning[69, 70].Cao and Liu introduced a novel method, ReLMKG, designed for the multi-hop Knowledge Base Question Answering (KBQA) task. The approach leverages both the implicit knowledge embedded in a pre-trained language model and the explicit knowledge from KGs. They transferred the KG into a textual format, aligning it with the question in a shared space. The textualized graph and the question are input for the pre-trained language model which acts as the question-path encoder. The lack of the inherent topological arrangement in the textual KG, which holds essential explicit knowledge for model reasoning, led to the creation of a path-centric Graph Neural Network (GNN) as the reasoning element. The instructions for the reasoning module are derived from the outputs of the pre-trained language model. Sen et al. introduced an innovative approach to address queries, employing a KGQA model to extract information from a KG. Simultaneously, a language model is employed to analyze the question and facts, enabling the derivation of a well-reasoned answer. Baek et al. utilized a similarity metric between KG facts and questions. This approach was employed to retrieve pertinent facts, subsequently enhancing the prompt of a language model.
@@
Query Generation from text: This portion of the paper specifically addresses Research Question 5, which examines the effectiveness of LLMs in generating natural language queries, particularly focusing on the conversion of textual descriptions into query languages like SPARQL or Cypher for KGs. Constructing SPARQL queries from natural language questions is challenging, as it demands an understanding of both the question itself and the underlying patterns within the KG [74]. Rony et al. introduced a novel embedding technique called SGPT, tailored to encode linguistic features from the question, and suggested training methods that make use of a pre-trained language model to produce a SPARQL query. Kovriguina et al. proposed a one-shot generative approach, termed SPARQLGEN, which aims to generate SPARQL queries by augmenting LLMs with pertinent context provided in a single prompt. This prompt encompasses the question, an RDF subgraph necessary for answering the question, and an example of a correct SPARQL query for a different question. Pliukhin et al. devised a method that incorporates diverse data sources in the SPARQL generation prompt, encompassing the question, an RDF subgraph essential for answering the question, and an illustration of a correct SPARQL query. Their method suggests a couple of enhancements to the SPARQLGEN approach.
@@
Querying LLMs with SPARQL: Saeed et al. explored the possibility of querying LLMs with SQl, initiating new directions for research exploration. While SQL is a powerful tool for structured datasets with defined schemas, it falls short in immediately handling information presented as unstructured text. The statement discusses an illustration of a Database-First (DB-first) approach that harnesses the capabilities of LLMs along with traditional Database Management Systems (DBMSs). This combination is utilized to establish a hybrid query execution environment, suggesting an integration of LLMs with conventional database systems for enhanced performance. This concept might also align with querying LLMs with SPARQL, which might obtain hidden relations in unstructured data.
@@
KG Chatbots: Omar et al. conducted a detailed examination and comparison between conversational AI models like ChatGPT and traditional Question-Answering Systems (QASs) designed for KGs and suggested merging the attributes of both with the goal of creating KG Chatbots.
5 Statistics Analysis and Open Challenges
@@
Statistics about used LLMs and KGs in Approaches: Our analysis categorizes research papers based on the use of LLMs and KGs. Figure 2 illustrates this, with the y axis representing the frequency of appearances and the x axis listing the specific KGs and LLMs. The data reveal that Freebase is the most commonly utilized KG in the reviewed literature. Regarding LLMs, Bert and GPT-3 emerge as the most frequently employed models. Notably, in the context of KG-enhanced LLMs, Bert is the only model that has been utilized.
Figure 2: LLMs and KGs used in papers
@@
Open Challenges: We still need research to find techniques to incorporate knowledge reliably into the answers of LLMs and to go for smaller-sized LLMs without loosing the capabilities of LLMs to reason on their input. One way seems to be to incorporate the knowledge from KGs reliably into the inference process of LLMs and to exclude the knowledge from the training data because these facts are then not needed anymore to be stored in the neural network of the LLMs [79] increasing its number of parameters. Running LLMs with fewer parameters reduces the energy needed and hence the carbon footprint of running LLMs could be drastically decreased. Hence we see a demand for a better separation of knowledge (to be managed in KGs in a condensed and reliable way) and natural language understanding (based on a well-chosen minimal training data set of high quality not addressing the knowledge of the given KG). How to retrieve this well-chosen minimal training data set is a non-trivial question and may open up a new research direction itself. With a clear separation of knowledge and natural language understanding, fine-tuning to specific domains may also become obsolete or at least can be reduced to a minimum effort. New applications of LLMs like Personal KG-enhanced LLMs, which can imitate the style of writing of each individual by fine-tuning from email and chat conversations and based on a Personal KG containing the (private) knowledge of the individual might become true.
@@
Besides reducing the complexity of LLMs there might be another research direction on more complex architectures mimicking the working of the human brain even more towards the holy grail of artificial intelligence, the artificial general intelligence (AGI), to accomplish any intellectual task that human beings or animals can perform [80]. In these architectures, LLMs or advanced variants might play only a minor role by being responsible for processing the input and/or verbalizing the conclusion of the overall process. KGs might be still a standard way for administrating the knowledge in a reliable and space-efficient way in these architectures for AGI.
6 Conclusion
@@
In conclusion, this comprehensive survey has significantly contributed to the understanding and advancement of the integration between LLMs and KGs. Through our meticulous exploration across various sections, we have addressed critical research questions, shedding light on how LLMs can enhance KGs in areas like generating descriptive textual information, ontology generation, inconsistency detection, fact-checking, natural language query generation, and KG Question Answering systems. Our study not only delves into the conventional applications but also pioneers in exploring under-discussed areas like KG validation and multi-hop question generation and answering.

### v2

@@ 4.1.1. Multi-Hop Question Generation:
The goal of multi-hop question generation is to formulate questions demanding advanced reasoning over multiple sentences, including their associated answers (Li et al., 2023a). Li et al. created a KG Enhanced Language Model (KGEL), aiming to replicate human-like reasoning for multi-hop questions. Their method for generating multi-hop questions reflects the way humans approach complex questioning. Their strategy encompasses three key phases: (1) Gaining an understanding of the context by employing a pre-trained GPT-2 language model, (2) Merging information and reasoning through a sophisticated KG and an answer-aware dynamic Graph Attention Network (GAT), and (3) Executing the question generation using a multi-head attention generation module that capitalizes on improved latent representations. Aigo et al. research primary emphasis revolved around formulating inquiries by employing KGs in conjunction with the T5 language model. They harnessed the language model’s capabilities to generate questions while also applying a technique to obstruct self-attention in the encoder, all to train the model with a deliberate emphasis on maintaining the inherent structure of the graph. They harnessed the language model’s capabilities to generate questions while also applying a technique to obstruct self-attention in the encoder, all to train the model with a deliberate emphasis on maintaining the inherent structure of the graph. Their research didn’t target multi-hop question generation.
4.1.2. Complex or Multi-hop Question Answering:
@@
Complex question answering over KGs involves multiple subjects, conveys interconnected relations, includes numerical operations (Lan et al., 2021), and demands multi-hop reasoning (Yang et al., 2018; Etezadi and Shamsfard, 2023). Cao and Liu introduced a novel method, ReLMKG, designed for the multi-hop Knowledge Base Question Answering (KBQA) task. The approach leverages both the implicit knowledge embedded in a pre-trained language model and the explicit knowledge from KGs. They transferred the KG into a textual format, aligning it with the question in a shared space. The textualized graph and the question are input for the pre-trained language model, which acts as the question-path encoder. The lack of the inherent topological arrangement in the textual KG, which holds essential explicit knowledge for model reasoning, led to the creation of a path-centric Graph Neural Network (GNN) as the reasoning element. The instructions for the reasoning module are derived from the outputs of the pre-trained language model. Sen et al. introduced an innovative approach to address queries, employing a KGQA model to extract information from a KG. Simultaneously, a language model is employed to analyze the question and facts, enabling the derivation of a well-reasoned answer. Baek et al. utilized a similarity metric between KG facts and questions. This approach was employed to retrieve pertinent facts, subsequently enhancing the prompt of a language model.
4.1.3. Query Generation from text:
@@
This portion of the paper specifically addresses Research Question 6, which examines the effectiveness of LLMs in generating natural language queries. It particularly focuses on converting textual descriptions into query languages like SPARQL or Cypher for KGs. Constructing SPARQL queries from natural language questions is challenging, as it demands an understanding of both the question itself and the underlying patterns within the KG (Rony et al., 2022). Rony et al. introduced a novel embedding technique called SGPT, tailored to encode linguistic features from the question, and suggested training methods that make use of a pre-trained language model to produce a SPARQL query. Kovriguina et al. proposed a one-shot generative approach, termed SPARQLGEN, which aims to generate SPARQL queries by augmenting LLMs with pertinent context provided in a single prompt. This prompt encompasses the question, an RDF subgraph necessary for answering the question, and an example of a correct SPARQL query for a different question. Pliukhin et al. devised a method that incorporates diverse data sources in the SPARQL generation prompt, encompassing the question, an RDF subgraph essential for answering the question, and an illustration of a correct SPARQL query. Their method suggests a couple of enhancements to the SPARQLGEN approach.
4.1.4. Querying LLMs with SPARQL:
@@
Saeed et al. explored the possibility of querying LLMs with SQl, initiating new directions for research exploration. While SQL is a powerful tool for structured datasets with defined schemas, it falls short in immediately handling information presented as unstructured text. The statement discusses an illustration of a Database-First (DB-first) approach that harnesses the capabilities of LLMs along with traditional Database Management Systems (DBMSs). This combination is utilized to establish a hybrid query execution environment, suggesting an integration of LLMs with conventional database systems for enhanced performance. This concept might also align with querying LLMs with SPARQL, which might obtain hidden relations in unstructured data.
4.1.5. KG Chatbots:
@@
Omar et al. conducted a detailed examination and comparison between conversational AI models like ChatGPT and traditional Question-Answering Systems (QASs) designed for KGs and suggested merging the attributes of both with the goal of creating KG Chatbots.
5. Statistics Analysis and Open Challenges
5.1. Statistics about used LLMs and KGs in Approaches
@@
Our analysis categorizes research papers based on the use of LLMs and KGs. Figure 2 illustrates this, with the y-axis representing the frequency of appearances and the x-axis listing the specific KGs and LLMs. The data reveal that Freebase is the most commonly utilized KG in the reviewed literature. Regarding LLMs, Bert and GPT-3 emerge as the most frequently employed models.
5.2. Open Challenges
@@
We still need research to find techniques to incorporate knowledge reliably into the answers of LLMs and to go for smaller-sized LLMs without losing the capabilities of LLMs to reason on their input. One way seems to be to incorporate the knowledge from KGs reliably into the inference process of LLMs and to exclude the knowledge from the training data because these facts are then not needed anymore to be stored in the neural network of the LLMs (Meng et al., 2022) increasing its number of parameters. Also, incorporating only code of a core fragment of a programming language or a query language like SPARQL (Groppe et al., 2009a) and XPath (Groppe et al., 2006a; Groppe and Groppe, 2006b) without redundant language constructs in the training data may reduce the number of needed parameters in LLMs. To reduce the number of needed parameters, future work may also consider generating code and queries in a limited set of programming and query languages and transforming the generated code to the requested programming and query languages. For example, automated transformations between XML and Semantic Web data format and query languages are well known (Droop et al., 2007, 2008a, 2008b; Groppe et al., 2008, 2011), and also reformulating queries based on a given transformation of the data (Groppe et al., 2006b). Other techniques may consider only those queries in the training data, which can return a result and are hence satisfiable (Groppe and Groppe, 2006a, 2008; Groppe et al., 2009b; Groppe and Groppe, 2006b). Running LLMs with fewer parameters reduces the energy needed and, hence, the carbon footprint of running LLMs. We see a demand for a better separation of knowledge (to be managed in KGs in a condensed and reliable way) and natural language understanding (based on a well-chosen minimal training data set of high quality not addressing the knowledge of the given KG). How to retrieve this well-chosen minimal training data set is a non-trivial question and may open up a new research direction itself. With a clear separation of knowledge and natural language understanding, fine-tuning to specific domains may also become obsolete or at least can be reduced to a minimum effort. New applications of LLMs like Personal KG-enhanced LLMs, which can imitate the style of writing of each individual by fine-tuning from email and chat conversations and based on a Personal KG containing the (private) knowledge of the individual might become true.
@@
Besides reducing the complexity of LLMs there might be another research direction on more complex architectures mimicking the working of the human brain even more towards the holy grail of artificial intelligence, the artificial general intelligence (AGI), to accomplish any intellectual task that human beings or animals can perform (Shevlin et al., 2019). In these architectures, LLMs or advanced variants might play only a minor role by being responsible for processing the input and/or verbalizing the conclusion of the overall process. KGs might still be a standard way for administrating knowledge in a reliable and space-efficient way in these architectures for AGI.
6. Conclusion
@@
In conclusion, this comprehensive survey has significantly contributed to the understanding and advancement of the integration between LLMs and KGs. Through our meticulous exploration across various sections, we have addressed critical research questions, shedding light on how LLMs can enhance KGs in areas such as generating descriptive textual information, ontology generation, inconsistency detection, fact-checking, KG completion, and KG embedding. Our study also investigates how KGs can improve LLMs and explores the cooperation between LLMs and KGs. Specifically, it examines under-discussed areas in KG question answering, including query generation from text and multi-hop question generation and answering.
Acknowledgements.
The German Research Foundation funds this work under project number 490998901.